# Simulating and investigating various dynamic aspects of $\rm{H}_{2}\rm{O}$-related hydrogen bond model

### v1

@@ Abstract
A simple $\rm{H}_{2}\rm{O}$-related hydrogen bond model, modified from the Jaynes–Cummings model, is proposed and its various dynamic aspects are investigated theoretically. In this model, the formation and breaking processes of hydrogen bond are accompanied by the creation and annihilation of the thermal phonon of the medium. A number of simplifying assumptions about the dynamics of the molecules involved are used. Rotating wave approximation is applied under consideration of the strong-coupling condition. Dissipative dynamics under the Markovian approximation is obtained through solving the quantum master equation — Lindbladian. The probabilities of reaction channels involving hydrogen bond depending on the parameters of the external environment, are obtained. Differences between unitary and dissipative evolutions are disciussed. Consideration is given to the effect of all kinds of potential interactions and dissipations on evolution. Consideration is also given to the reverse processes (inflows) of dissipations. The results show that the magnitude changes of the interactions and dissipations have slight effect on the formation of hydrogen bond, but the variation of the reverse processes of dissipations significantly affect the formation of hydrogen bond. According to the findings, the dynamics of $\rm{H}_{2}\rm{O}$-related hydrogen bond model can be controlled by selectively choosing system parameters. The results will be used as a basis to extend the research to more complex chemical and biological model in the future.
@@ II Hydrogen bond model
The hydrogen bond dynamics in media as a dynamics of polariton of the group of real particles — two molecules and quasiparticles (photons and phonons), is represented. Strictly speaking, the quantum state of polariton has the form $\lambda_{photon}|photon\rangle+\lambda_{phonon}|phonon\rangle+\lambda_{particles}|particles\rangle$. This is hardly possible to deal with this form, because any interacton: photon-phonon, phonon-atoms, photon-atoms represents nontrivial task. Thus, the initial photon can transform to a few phonons. So it is accepted to call the polariton of phonons if there is no explicit photons in advance. This terminology is followed and the system in the framework of Jaynes–Cummings scheme is presented with Hamiltonian $a^{{\dagger}}\sigma+a\sigma^{{\dagger}}+a^{{\dagger}}a+\sigma^{{\dagger}}\sigma$, where the field operator $a$ relates to our conditional phonon, and operator $\sigma$ — to real particles.
@@
$g(a^{\dagger}+a)\sigma^{\dagger}\sigma$ (1)
where $a$ is the annihilation operator of the phonon of the selected mode, $\sigma$ is the relaxation operator of the atom. This approach is extremely computationally expensive, since it requires explicit inclusion of phonons in the basic states of the model, which immediately causes a huge increase in the required memory and does not allow scaling the model to multi-molecular structures to study new collective effects. In addition, the case of atomic excitations is very different from Eq. (1), because their energy is several orders of magnitude greater than the energy of a single phonon, so that a term of the form Eq. (1) corresponds only to dephasing, but not to direct interaction of phonons with matter.
@@
where $\sigma_{hb}$ is the operator of hydrogen bond formation accompanied by phonon emission, and $a_{phn}$ is the operator of phonon absorption leading to bond decay. Direct inclusion of such a term in the Hamiltonian is also unacceptable for us, since explicit phonons again arise in the basic states. An approximation in the form of the average number of phonons $n_{av}$ at the resonance frequency sensitive to hydrogen bonding is applied; since the operators of phonon annihilation and creation are proportional to the square root of their number $\sqrt{n}$, $H_{int}$ is replaced with the operator
@@
How to relate $n_{av}$ to the temperature of the environment of interacting molecules? The thermally stable state of the phonon field inside the cavity will is introduced and has the following form
@@
where $hbar$ is reduced Planck constant, $\omega$ is phononic mode, $n$ is phonon number, $T$ is the cavity temperature at a given frequency mode $\omega$, $K$ is the Boltzmann constant, $c$ is normalization factor. $\gamma_{in}$ and $\gamma_{out}$, which are the intensities of the inflow and dissipation of phonons into and out of a notional cavity around the molecules, are proposed to form a geometric progression with the denominator [49]
@@
where $\gamma^{in}$ refers to the overall spontaneous inflow rate, $\gamma^{out}$ refers to the overall spontaneous emission rate. A stable temperature occurs only when $\gamma_{in}<\gamma_{out}$. Knowing these coefficients, or knowing the temperature $T$ directly, $n_{av}$ is obtained. In practice, the coefficients are found only by optimizing them using neural networks based on the experimental results. The temperature at a fixed photonic mode is determined in a similar way. Photonic modes relate to transformations of electron states, and phononic modes relate to proton oscillations. Their energy is approximately 2 orders of magnitude less.
@@ II.1 The target model
Figure 1: (online color) Hydrogen bond formation mechanism. Scale model of two $\rm{H}_{2}\rm{O}$ molecules is shown in panel (a). Mechanism of ion attraction between hydrogen and oxygen atoms is shown in the form of a Ball–and–stick model in panel (b). Brown circles (balls) represent oxygen atom, red circles (balls) represent hydrogen atom (proton) in excited state, black solid lines stand for covalent bond and gray dotted lines represent ion attraction; especially, black dashed line indicate stretched covalent bond and green ball in panel (a) represent a proton in ground state.
@@
Interaction between two water molecules causes micro-oscillation of the hydrogen atom in one of water molecules, it allows hydrogen bonds to form between water molecules, but this process does not break covalent bond. The hydrogen bond formation mechanism is as shown in Fig. 1. In panel (a), when two $\rm{H}_{2}\rm{O}$ molecules that are moving freely are far apart, hydrogen bond cannot form between the molecules. However, when the molecules move closer together, the hydrogen atom of one molecule and the oxygen atom of another molecule attract each other and stable hydrogen bond is obtained. In panel (b), the oxygen atom (proton) of one molecule “donor” is attracted by the oxygen atom of another molecule “receptor” to produce the tunneling effect. The proton tunneling in the normal state does not destroy the covalent bond between proton and its parent molecule, but only deforms it. In the case without considering of electrons, the effect of some stretching of the covalent bond on the hydrogen bond formation is ignored, and it is considered as a normal state with a covalent bond to avoid cluttering the target model. In the model with electrons, this difference is no longer ignored. Theoretically, in addition to 2 covalent bonds, for the oxygen atom, there are actually 2 more hydrogen bonds with protons from neighboring molecules. The oxygen atom and its covalent and hydrogen bonds form a tetrahedron.
@@
A hydrogen bond between water molecules occurs when they approach each other at a distance $|0\rangle_{dist}$, allowing one of the protons, covalently bonded to the parent molecule, to tunnel between it and another neighbouring molecule, while maintaining the existing covalent bond. The qubit $|d\rangle_{dist}$ describing the relative position of proton in the system, is introduced, and $d\in\{-1,0,1,2\}$. $dist=2$ means that the molecules are separated by a large distance, at which the formation of a hydrogen bond is impossible in principle. The state $dist=1$ means that the molecules are approaching each other at a critical distance, and although it does not allow the creation of a hydrogen bond, it allows the molecules to tunnel to the state $dist=0$, where the bond becomes possible. The state $dist=-1$ means the presence of a stable hydrogen bond, here it is an irreversible process. The influence of temperature on formation of hydrogen bond is investigated, because change of temperature will cause inflow of phonon. Thence, in this situation this process becomes reversible, that means the breaking of hydrogen bonds becomes possible.
@@
where $H_{i}^{basis}$ is diagonal in the standard basis, and $H_{i}^{cogerence}$ is not diagonal in this basis, then the term $H_{i}^{basis}$ is added to the total Hamiltonian in any case, and $H_{i}^{cogerence}$ is added only if the condition $A$ is satisfied. This remark about the conditional Hamiltonian will be valid for further models as well. This method achieves equality of rest energies for all terms of the general Hamiltonian. So the condition operator $Cond$ means only the inclusion of coherent terms in the Hamiltonian. For example, the proton (and, subsequently, the electron) jumps only if $d=0$, that is, when the molecules are close. But in this case, the stationary energy of both the proton and the electron is always present in the Hamiltonian; only at a large distance between the molecules is it as if “frozen”, and does not lead to a change in the state of the proton (and, subsequently, the electron), and as soon as the molecules are close, the motions of these particles begin. The standard basis thus turns out to be distinguished from all others. In this work, a single hydrogen bond between a pair of water molecules, which will be extended to more hydrogen bonds using Eq. (6) in the future, is considered.
@@ II.2 States and Hamiltonian
According to Eq. (7), when $n_{av}=1$, the Hilbert space of the $\rm{H}_{2}\rm{O}$-related hydrogen bond model will consist of the following 7 basic states, which are shown in Tab. 1. And the Ball-and-stick forms of these states are shown in Fig. 2.
@@
$H_{hyd}$ is represented by the Hamiltonian of a system of two water molecules connected by hydrogen bond and has the following form
$H_{hyd}=H_{dist}Cond(p=1)+H_{prot}Cond(d=0)$ (8)
@@
The complete quantum system ultimately constructed by the total Hamiltonian $H_{hyd}$ can be intuitively represented in the form of a network, which is shown in Fig. 3. Interactions, dissipations and inflows between states can be intuitively seen in this figure, and the energy wandering of the whole system is determined clearly.
@@ II.3 Quantum master equation
where $\gamma_{phn}$ is the dissipation intensity for the escape of phonons from cavity to external environment. Similarly, the introduction of two quasi-decoherence factors, which are similar to the form of $A_{phn},\ A_{phn}^{{\dagger}}$ and related to the qubit $|d\rangle_{dist}$ of the basic state, is as follows
@@
where $\gamma_{bond}$ is the dissipation intensity for the micromotions of molecules that must calm down to form a bond, and $\gamma_{isol}$ is the dissipation intensity for the macromotions of molecules in the medium. And
@@
where $\mathcal{L}=L^{in}+L^{out}$. $L^{in}$ describes the inflow process and $L^{out}$ describes the dissipation process. And
@@ III Numerical method
The third step is to scale the density matrix to maintain its positive definite, Hermitian matrix, unit trace and other properties.
@@ IV Simulations and results
Various dynamic aspects of $\rm{H}_{2}\rm{O}$-related hydrogen bond model are shown as below five subsections.
@@ IV.1 Comparison between unitary and dissipative evolutions
Dissipation intensities are taken into consideration to calculate the dissipative evolution of an open system with thermal phonon entering and exiting. The QME is numerically calculated using Eq. (20). In addition to initial state $|0\rangle|0\rangle|1\rangle$, the initial state $|1\rangle|1\rangle|0\rangle$ — critical state, is also considered. Critical state means the critical point of hydrogen bond breaking is reached. Comparison between these two initial states is carried out in Fig. 5. As shown in both panels, the periodic oscillations disappear, replaced by dissipations. The time of hydrogen bond formation is about $1\rm{ps}$ ($10^{-12}s$) before the probability reaches a stable level. In panel (a), the system tends to form a stable hydrogen bond when the initial state is $|0\rangle|0\rangle|1\rangle$. On the contrary, in panel (b), the system tends to break the hydrogen bond when the initial state is $|1\rangle|1\rangle|0\rangle$. This is because the state $|0\rangle|0\rangle|1\rangle$ is stretched and ground, and the proton of donor at this time is at lower energy lever. Thus, it is more inclined to form a stable hydrogen bond. The state $|1\rangle|1\rangle|0\rangle$ means that the critical point of hydrogen bond formation and breaking is obtained, and the hydrogen bond breaks more easily to release two free water molecules.
@@ IV.2 The effect of interactions on evolution
In the first row of Fig. 6, when the initial state is $|0\rangle|0\rangle|1\rangle$, result is found that the larger interaction strengths $g_{dist}$ and $g_{prot}$, the lower the probability of forming stable hydrogen bond. However, in the second row, when initial state is $|1\rangle|1\rangle|0\rangle$, the opposite result is found. In summary, in the case of $|0\rangle|0\rangle|1\rangle$, the interaction strengths have the negative effect on the formation of stable hydrogen bond — hindering the hydrogen bond formation; in the case of $|1\rangle|1\rangle|0\rangle$, the interaction strengths have the positive effect on the formation of stable hydrogen bond — promoting the hydrogen bond formation. Besides, when the initial state is $|0\rangle|0\rangle|1\rangle$, the probabilities on the heat maps are all on the range of [0.5, 1]0.51[0.5,\ 1][ 0.5 , 1 ]; When the initial state is $|1\rangle|1\rangle|0\rangle$, the probabilities on the heat maps are all on the range of [0, 0.5)00.5[0,\ 0.5)[ 0 , 0.5 ).
@@
In addition to the above results, the effect of different interactions strengths on the time required for the system to reach stable is also obtained: the smaller interactions strengths, the more time required to reach stable. When $g_{dist}=g_{prot}=0.5g$ system requires the most time to reach stable.

### v2

@@ Abstract
A simple $\rm{H}_{2}\rm{O}$-related hydrogen bond model, modified from the Jaynes–Cummings model, is proposed and its various dynamic aspects are investigated theoretically. In this model, the formation and breaking processes of hydrogen bond are accompanied by the creation and annihilation of the thermal phonon of the medium. A number of simplifying assumptions about the dynamics of the molecules involved are used. Rotating wave approximation is applied under consideration of the strong-coupling condition. Dissipative dynamics under the Markovian approximation is obtained through solving the quantum master equation — Lindbladian. The probabilities of reaction channels involving hydrogen bond depending on the parameters of the external environment, are obtained. Differences between unitary and dissipative evolutions are discussed. Consideration is given to the effect of all kinds of potential interactions and dissipations on evolution. Consideration is also given to the reverse processes (inflows) of dissipations. The results show that the magnitude changes of the interactions and dissipations have slight effect on the formation of hydrogen bond, but the variation of the inflows significantly affects the formation of hydrogen bond. According to the findings, the dynamics of $\rm{H}_{2}\rm{O}$-related hydrogen bond model can be controlled by selectively choosing system parameters. The results will be used as a basis to extend the research to more complex chemical and biological model in the future.
@@ II Hydrogen bond model
The hydrogen bond dynamics in media as a dynamics of polariton of the group of real particles — two molecules and quasiparticles (photons and phonons), is represented. Strictly speaking, the quantum state of polariton has the form
$\lambda_{photon}|photon\rangle+\lambda_{phonon}|phonon\rangle+\lambda_{particle}|particle\rangle$ (1)
This is hardly possible to deal with this form, because any interaction (photon–phonon, phonon–atom, photon–atom) represents nontrivial task. Thus, the initial photon can transform to a few phonons. So it is accepted to call the polariton of phonons if there is no explicit photons in advance. This terminology is followed and the system in the framework of Jaynes–Cummings scheme is presented with Hamiltonian
$a^{{\dagger}}\sigma+a\sigma^{{\dagger}}+a^{{\dagger}}a+\sigma^{{\dagger}}\sigma$ (2)
where the field operator $a$ relates to our conditional phonon, and operator $\sigma$ — to real particles.
@@
$g(a_{\omega}^{\dagger}+a_{\omega})\sigma_{\omega}^{\dagger}\sigma_{\omega}$ (3)
where $a_{\omega}$ is the annihilation operator of the phonon of the selected mode $\omega$, $\sigma_{\omega}$ is the relaxation operator of the atom. This approach is extremely computationally expensive, since it requires explicit inclusion of phonons in the basic states of the model, which immediately causes a huge increase in the required memory and does not allow scaling the model to multi-molecular structures to study new collective effects. In addition, the case of atomic excitations is very different from Eq. (3), because their energy is several orders of magnitude greater than the energy of a single phonon, so that a term of the form Eq. (3) corresponds only to dephasing, but not to direct interaction of phonons with matter.
@@
where $\sigma_{hb}$ is the operator of hydrogen bond formation accompanied by phonon emission, and $a_{phn}$ is the operator of phonon absorption leading to bond decay. Direct inclusion of such a term in the Hamiltonian is also unacceptable for us, since explicit phonons again arise in the basic states. An approximation in the form of the average number of phonons $n_{av}$ at the resonance frequency sensitive to hydrogen bonding is applied; since the operators of phonon annihilation and creation are proportional to the square root of their number $\sqrt{n_{av}}$, $H_{int}$ is replaced with the operator
@@
How to relate $n_{av}$ to the temperature of the environment of interacting molecules? The thermally stable state of the phonon field inside the cavity is introduced and has the following form
@@
where $\hbar$ is reduced Planck constant, $\omega$ is phononic mode, $n$ is phonon number, $T$ is the cavity temperature at a given frequency mode $\omega$, $K$ is the Boltzmann constant, $c$ is normalization factor. $\gamma_{in}$ and $\gamma_{out}$, which are the intensities of the inflow and dissipation of phonons into and out of a notional cavity around the molecules, are proposed to form a geometric progression with the denominator [49]
@@
where $\gamma^{in}$ refers to the overall spontaneous inflow rate, $\gamma^{out}$ refers to the overall spontaneous emission rate. A stable temperature occurs only when $\gamma_{in}<\gamma_{out}$. Knowing these coefficients, or knowing the temperature $T$ directly, $n_{av}$ is obtained. In practice, the coefficients are found only by optimizing them using neural networks based on the experimental results. The temperature at a fixed photonic mode is determined in a similar way. Photonic modes relate to transformations of electron states, and phononic modes relate to proton oscillations.
@@ II.1 The target model
Figure 1: (online color) Hydrogen bond formation mechanism. Scale model of two $\rm{H}_{2}\rm{O}$ molecules is shown in panel (a). Mechanism of ion attraction between hydrogen and oxygen atoms is shown in the form of a ball–and–stick model in panel (b). Brown circles (balls) represent oxygen atom, red circles (balls) represent hydrogen atom in excited state, black solid lines stand for covalent bond and gray dotted lines represent ion attraction; especially, black dashed line indicate stretched covalent bond and green ball in panel (a) represent a proton in ground state.
@@
Interaction between two water molecules causes micro-oscillations of the hydrogen atom in one of water molecules, it allows hydrogen bond to form between water molecules, but this process does not break covalent bond. The hydrogen bond formation mechanism is as shown in Fig. 1. In panel (a), when two $\rm{H}_{2}\rm{O}$ molecules that are moving freely are far apart, hydrogen bond cannot form between the molecules. However, when the molecules move closer together, the hydrogen atom (proton) of one molecule and the oxygen atom of another molecule attract each other and stable hydrogen bond is obtained. In panel (b), the hydrogen atom of one molecule “donor” is attracted by the oxygen atom of another molecule “receptor” to produce the tunneling effect. The proton tunneling in the normal state does not destroy the covalent bond between proton and its parent molecule, but only deforms it. In the case without considering of electrons, the effect of some stretching of the covalent bond on the hydrogen bond formation is ignored, and it is considered as a normal state with a covalent bond to avoid cluttering the target model. In the model with electrons, this difference is no longer ignored. Theoretically, in addition to 2 covalent bonds, for the oxygen atom, there are actually 2 more hydrogen bonds with protons from neighboring molecules. The oxygen atom and its covalent and hydrogen bonds form a tetrahedron.
@@
A hydrogen bond between water molecules occurs when they approach each other at a distance $|0\rangle_{dist}$, allowing one of the protons, covalently bonded to the parent molecule, to tunnel between it and another neighbouring molecule, while maintaining the existing covalent bond. The qubit $|d\rangle_{dist}$ describing the relative position of proton in the system, is introduced, and $d\in\{-1,0,1,2\}$. $dist=2$ means that the molecules are separated by a large distance, at which the formation of a hydrogen bond is impossible in principle. The state $dist=1$ means that the molecules are approaching each other at a critical distance, and although it does not allow the creation of a hydrogen bond, it allows the molecules to tunnel to the state $dist=0$, where the bond becomes possible. The state $dist=-1$ means the presence of a stable hydrogen bond, here it is an irreversible process. However, the influence of temperature on formation of hydrogen bond is investigated, because change of temperature will cause inflow of phonon. Thence, in this situation this process becomes reversible, that means the breaking of hydrogen bonds becomes possible.
@@
where $H_{i}^{basis}$ is diagonal in the standard basis, and $H_{i}^{cogerence}$ is not diagonal in this basis. The term $H_{i}^{basis}$ is added to the total Hamiltonian in any case, and $H_{i}^{cogerence}$ is added only if the condition $A$ is satisfied. This remark about the conditional Hamiltonian will be valid for further models as well. This method achieves equality of rest energies for all terms of the general Hamiltonian. So the condition operator $Cond$ means only the inclusion of coherent terms in the Hamiltonian. For example, the proton (and, subsequently, the electron) jumps only if $d=0$, that is, when the molecules are close. Only at a large distance between the molecules is it as if “frozen”, and does not lead to a change in the state of the proton (and, subsequently, the electron). And as soon as the molecules are close, the motions of these particles begin. The standard basis thus turns out to be distinguished from all others. In this work, a single hydrogen bond between a pair of water molecules, which will be extended to more hydrogen bonds using Eq. (8) in the future, is considered.
@@ II.2 States and Hamiltonian
According to Eq. (9), when $n_{av}=1$, the Hilbert space of the $\rm{H}_{2}\rm{O}$-related hydrogen bond model will consist of the following 7 basic states, which are shown in Tab. 1. And the ball–and–stick forms of these states are shown in Fig. 2.
@@
$H_{hb}$ is represented by the Hamiltonian of a system of two water molecules connected by hydrogen bond and has the following form
$H_{hb}=H_{dist}Cond(p=1)+H_{prot}Cond(d=0)$ (10)
@@
The complete quantum system ultimately constructed by the total Hamiltonian $H_{hb}$ can be intuitively represented in the form of a network, which is shown in Fig. 3. Interactions, dissipations and inflows between states can be intuitively seen in this figure, and the energy wandering of the whole system is determined clearly.
@@ II.3 Quantum master equation
And $\gamma_{phn}$ is the dissipation intensity for the escape of phonons from cavity to external environment. Similarly, the two quasi-decoherence factors, which are related to the qubit $|d\rangle_{dist}$ of the basic state, are defined as follows
@@
And $\gamma_{bond}$ is the dissipation intensity for the micromotions of molecules that must calm down to form a bond, $\gamma_{isol}$ is the dissipation intensity for the macromotions of molecules in the medium.
@@
where $\mathcal{L}=L^{in}+L^{out}$. $L^{in}$ describes the inflow process and $L^{out}$ describes the dissipation process. Then
@@ III Numerical method
The third step is to scale the density matrix to maintain its positive definiteness, Hermitian, normalization and other properties.
@@ IV Simulations and results
Various dynamic aspects of $\rm{H}_{2}\rm{O}$-related hydrogen bond model are shown as below five subsections. The supercomputer resources [58] is used to obtain the results of simulations.
@@ IV.1 Comparison between unitary and dissipative evolutions
Dissipation intensities are taken into consideration to calculate the dissipative evolution of an open system with thermal phonon entering and exiting. The QME is numerically calculated using Eq. (22). In addition to initial state $|0\rangle|0\rangle|1\rangle$, the initial state $|1\rangle|1\rangle|0\rangle$ — critical state, is also considered. Critical state means the critical point of hydrogen bond breaking is reached. Comparison between these two initial states is carried out in Fig. 5. As shown in both panels, the periodic oscillations disappear, replaced by dissipations. The time of hydrogen bond formation is about $1\rm{ps}$ ($10^{-12}s$) before the probability reaches a stable level. In panel (a), $\rho_{|-1\rangle|0\rangle|0\rangle}>\rho_{|2\rangle|1\rangle|0\rangle}$, thus the system tends to form a stable hydrogen bond when the initial state is $|0\rangle|0\rangle|1\rangle$. On the contrary, in panel (b), $\rho_{|-1\rangle|0\rangle|0\rangle}<\rho_{|2\rangle|1\rangle|0\rangle}$, the system tends to break the hydrogen bond when the initial state is $|1\rangle|1\rangle|0\rangle$. This is because the state $|0\rangle|0\rangle|1\rangle$ is stretched and ground, and the proton of donor at this time is at lower energy lever. Thus, it is more inclined to form a stable hydrogen bond. The state $|1\rangle|1\rangle|0\rangle$ means that the critical point of hydrogen bond formation and breaking is obtained, and the hydrogen bond breaks more easily to release two free water molecules.
@@ IV.2 The effect of interactions on evolution
In the first row of Fig. 6, when the initial state is $|0\rangle|0\rangle|1\rangle$, result is found that the larger interaction strengths $g_{dist}$ and $g_{prot}$, the lower the probability of stable hydrogen bond formation. However, in the second row, when initial state is $|1\rangle|1\rangle|0\rangle$, the opposite result is found. In summary, in the case of $|0\rangle|0\rangle|1\rangle$, the interaction strengths have the negative effect on the formation of stable hydrogen bond — hindering the hydrogen bond formation; in the case of $|1\rangle|1\rangle|0\rangle$, the interaction strengths have the positive effect on the formation of stable hydrogen bond — promoting the hydrogen bond formation. Besides, when the initial state is $|0\rangle|0\rangle|1\rangle$, the probabilities on the heat maps are all on the range of [0.5, 1]0.51[0.5,\ 1][ 0.5 , 1 ]; when the initial state is $|1\rangle|1\rangle|0\rangle$, the probabilities on the heat maps are all on the range of [0, 0.5)00.5[0,\ 0.5)[ 0 , 0.5 ). This is because according to the results in Fig. 5 (a), when the initial state is $|0\rangle|0\rangle|1\rangle$, the probability of the $|\rm{O}\cdots\rm{H}\rangle$ is greater than $0.8$, and in the first row of Fig. 6 the interaction strengths have slight effect on the evolution, so the probabilities represented by the points on the first row of the heat maps are almost all around $0.8$. Similarly, according to the Fig. 5 (b), when the initial state is $|1\rangle|1\rangle|0\rangle$, the probability of the $|\rm{O}\cdots\rm{H}\rangle$ is less than $0.2$, and in the second row of Fig. 6 the interaction strengths have also slight effect on the evolution, so the probabilities represented by the points on the second row of the heat maps are almost all around $0.2$.
@@
In addition to the above results, the effect of different interactions strengths on the time required for the system to reach stable is also obtained: the smaller interactions strengths, the more time required to reach stable.